# MODULI: Unlocking Preference Generalization via Diffusion Models for Offline Multi-Objective Reinforcement Learning

### v1

@@ Abstract
Multi-objective Reinforcement Learning (MORL) seeks to develop policies that simultaneously optimize multiple conflicting objectives, but it requires extensive online interactions. Offline MORL provides a promising solution by training on pre-collected datasets to generalize to any preference upon deployment. However, real-world offline datasets are often conservatively and narrowly distributed, failing to comprehensively cover preferences, leading to the emergence of out-of-distribution (OOD) preference areas. Existing offline MORL algorithms exhibit poor generalization to OOD preferences, resulting in policies that do not align with preferences. Leveraging the excellent expressive and generalization capabilities of diffusion models, we propose MODULI (Multi Objective DiffUsion planner with sLIding guidance), which employs a preference-conditioned diffusion model as a planner to generate trajectories that align with various preferences and derive action for decision-making. To achieve accurate generation, MODULI introduces two return normalization methods under diverse preferences for refining guidance. To further enhance generalization to OOD preferences, MODULI proposes a novel sliding guidance mechanism, which involves training an additional slider adapter to capture the direction of preference changes. Incorporating the slider, it transitions from in-distribution (ID) preferences to generating OOD preferences, patching, and extending the incomplete Pareto front. Extensive experiments on the D4MORL benchmark demonstrate that our algorithm outperforms state-of-the-art Offline MORL baselines, exhibiting excellent generalization to OOD preferences.
@@ 1 Introduction
Real-world decision-making tasks, such as robotic control [47, 29], autonomous driving [26], and industrial control [4], often require the optimization of multiple competing objectives simultaneously. It necessitates the trade-offs among different objectives to meet diverse preferences [39]. For instance, in robotic locomotion tasks, users typically focus on the robot’s movement speed and energy consumption [47]. If the user has a high preference for speed, the agent will move quickly regardless of energy consumption; if the user aims to save energy, the agent should adjust to a lower speed. One effective approach is Multi-objective Reinforcement Learning (MORL), which enables agents to interact with vector-rewarded environments and learn policies that satisfy multiple preferences [17, 38, 48]. These methods either construct a set of policies to approximate the Pareto front of optimal solutions [2, 13] or learn a preference-conditioned policy to adapt to any situation [33, 21]. However, exploring and optimizing policies for each preference while alleviating conflicts among objectives, requires extensive online interactions [3], thereby posing practical challenges due to high costs and potential safety risks.
@@
Offline MORL [50] proposes a promising solution that can learn preference-conditioned policies from pre-collected datasets with various preferences, improving data efficiency and minimizing interactions when deploying in high-risk environments. Most Offline MORL approaches focus on extending return-conditioned methods [43, 50] or encouraging consistent preferences through policy regularization [28]. However, real-world offline datasets often come from users with different behavior policies, making it difficult to cover the full range of preferences and leading to missing preference regions, i.e., OOD preferences, resulting in a shattered or narrow data distribution. As shown in fig. 1, we visualize the trajectory returns for Hopper-Amateur dataset. The Complete version of the dataset can comprehensively cover a wide range of preferences. In many cases, suboptimal policies can lead to Shattered or Narrow datasets, resulting in OOD preference regions as marked by the red circles. We then visualize the approximate Pareto front for MORvS baseline [50] and our MODULI in such a shattered dataset. We find that the current Offline MORL baselines exhibit poor generalization performance when learning policies from incomplete datasets. They fail to adapt to OOD preference policies. The approximate Pareto front of MORvS shows significant gaps in the OOD preference areas, indicating that the derived policies and preferences are misaligned. Therefore, we aim to derive a general multi-objective policy that can model a wide range of preference conditions with sufficient expressiveness and strong generalization to OOD preferences. Motivated by the remarkable capabilities of diffusion models in expressiveness and generalization, we propose MODULI (Multi Objective DiffUsion planner with sLIding guidance), transforming the problem of Offline MORL into conditional generative planning. In fig. 1 (right), MODULI demonstrated a better approximation of the Pareto front, achieving returns beyond the dataset. MODULI also patched the missing preference regions and exhibited excellent generalization for the OOD preferences.
@@ 2 Related Work
Most MORL research primarily focuses on online settings, aiming to train a general preference-conditioned policy [21, 32] or a set of single-objective policies [35, 47] that can generalize to arbitrary preferences. This paper considers an offline MORL setting that learns policy from static datasets without online interactions. PEDI [45] combines the dual gradient ascent with pessimism to multi-objective tabular MDPs. MO-SPIBB [43] achieves safe policy improvement within a predefined set of policies under constraints. These methods often require a priori knowledge about target preferences. PEDA [50] first develops D4MORL benchmarks for scalable Offline MORL with a general preference for high-dimensional MDPs with continuous states and actions. It also attempts to solve the Offline MORL via supervised learning, expanding return-conditioned sequential modeling to the MORL, such as MORvS [12], and MODT [6]. [28] integrates policy-regularized into the MORL methods to alleviate the preference-inconsistent demonstration problem. This paper investigates the generalization problem of OOD preference, aiming to achieve better generalization through more expressive diffusion planners and the guidance of transition directions by a slider adapter.
@@
Diffusion models (DMs) [19, 42] have become a mainstream class of generative models. Their remarkable capabilities in complex distribution modeling demonstrate outstanding performance across various domains [24, 40], inspiring research to apply DMs to decision-making tasks [7, 49, 46]. Applying DMs for decision-making tasks can mainly be divided into three categories [51]: generating long horizon trajectories like planner [22, 9, 1], serving as more expressive multimodal polices [44, 16] and performing as the synthesizers for data augmentation [31]. Diffusion models exhibit excellent generalization and distribution matching abilities [23, 27], and this property has also been applied to various types of decision-making tasks such as multi-task RL [18], meta RL [36] and aligning human feedback [10]. In this paper, we further utilize sliding guidance to better stimulate the generalization capability of DMs, rather than merely fitting the distribution by data memorization [15].
@@ 3 Preliminaries
The MORL problem can be formulated as a multi-objective Markov decision process (MOMDP) [5]. A MOMDP with $n$ objectives can be represented by a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\Omega,\gamma\rangle$, where $\mathcal{S}$ and $\mathcal{A}$ demote the state and action spaces, $\mathcal{P}$ is the transition function, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{n}$ is the reward function that outputs $n$-dim vector rewards $\bm{r}=\mathcal{R}(\bm{s},\bm{a})$, $\Omega$ is the preference space containing vector preferences $\bm{\omega}$, and $\gamma$ is a discount factor. The policy $\pi$ is evaluated for $m$ distinct preferences $\{\bm{\omega}_{p}\}_{p=1}^{m}$, resulting policy set be represented as $\{\pi_{p}\}_{p=1}^{m}$, where $\pi_{p}=\pi(\bm{a}|\bm{s},\bm{\omega}_{p})$, and $\bm{G}^{\pi_{p}}=\mathbb{E}_{\bm{a}_{t}\sim\pi_{p}}\left[\sum_{t}\gamma^{t}\mathcal{R}(\bm{s}_{t},\bm{a}_{t})\right]$ is the corresponding unweighted expected return. We say the solution $\bm{G}^{\pi_{p}}$ is dominated by $\bm{G}^{\pi_{q}}$ when $G_{i}^{\pi_{p}}<G_{i}^{\pi_{q}}$ for $\forall i\in[1,2,\cdots,n]$. The Pareto front $P$ of the policy $\pi$ contains all its solutions that are not dominated. In MORL, we aim to define a policy such that its empirical Pareto front is a good approximation of the true one. While not knowing the true front, we can define a set of metrics for relative comparisons among algorithms, e.g., hypervolume, sparsity, and return deviation, to which we give a further explanation in Section 5.
@@
where $\bm{w}^{t}\in\mathbb{R}^{D}$ is the standard Wiener process, and $f(t)=\frac{{\rm d}\log\alpha_{t}}{{\rm d}t}$, $g^{2}(t)=\frac{{\rm d}\sigma^{2}_{t}}{{\rm d}t}-2\sigma^{2}_{t}\frac{{\rm d}\log\alpha_{t}}{{\rm d}t}$. The SDE forward process in Equation 1 has an probability flow ODE (PF-ODE) reverse process from time $T$ to 00 [42]:
@@
Since $\bm{\epsilon}_{\theta}(\bm{x}^{t},t)$ can be considered as a predicted Gaussian noise added to $\bm{x}^{t}$, it is called the noise prediction model. With a well-train noise prediction model, DDIM [41] discretizes the PF-ODE to the first order for solving, i.e., for each sampling step:
@@
holds for any $0\leq t<s\leq T$. To enable the diffusion model to generate trajectories based on specified properties, Classifier Guidance (CG) [8] and Classifier-free Guidance (CFG) [20] are the two main techniques. We use CFG due to its stable and easy-to-train properties. CFG directly uses a conditional noise predictor to guide the solver:
@@ 4.1 Offline MORL as Conditional Generative Planning
We derive policies on the Offline MORL dataset $\mathcal{D}$ with trajectories $\tau=(\bm{\omega},\bm{s}_{0},\bm{a}_{0},\bm{r}_{0},\cdots,\bm{s}_{T-1},\bm{a}_{T-1},\bm{r}_{T-1})$ to respond to arbitrary target preferences $\bm{\omega}\in\Omega$ during the deployment phase. Previous Offline MORL [50] algorithms generally face poor expressiveness, unable to adapt to a wide range of preferences through a general preference-conditioned policy, resulting in sub-optimal approximate Pareto front. Inspired by the excellent expressiveness of diffusion models in decision making [1, 10], MODULI introduces diffusion models to model the policy. We define the sequential decision-making problem as a conditional generative objective $\max_{\theta}\mathbb{E}_{\tau\sim\mathcal{D}}\left[\log p_{\theta}\left(\bm{x}^{0}(\tau)\mid\bm{y}(\tau)\right)\right]$, where $\theta$ is the trainable parameter. The diffusion planning in MORL aims to generate sequential states $\bm{x}^{0}(\tau)=[\bm{s}_{0},\cdots,\bm{s}_{H-1}]$ with horizon $H$ to satisfy the denoising conditions $\bm{y}=[\bm{\omega},\bm{g}]$, where $\bm{\omega}$ is a specific preference. Similar to [50], $\bm{g}_{t}$ is defined as the vector-valued Returns-to-Go (RTG) for a single step, i.e. $\bm{g}_{t}=\sum_{t}^{T}\mathcal{R}(\bm{s}_{t},\bm{a}_{t})$. The trajectory RTG $\bm{g}$ is the average of the RTGs at each step.
@@
Next, we can sample batches of $\bm{x}^{0}(\tau)$ and $\bm{y}$ from the dataset and update $p_{\theta}$ as the modified score matching objective in eq. 3:
$\mathcal{L}(\theta)=\mathbb{E}_{(\bm{x}^{0},\bm{\omega},\bm{g})\sim\mathcal{D},t\sim U(0,T),\epsilon\sim\mathcal{N}(\bm{0},\bm{I})}\left[\|\bm{\epsilon}_{\theta}(\bm{x}^{t},t,\bm{\omega},\bm{g})-\bm{\epsilon}\|^{2}_{2}\right].$ (6)
MODULI uses the CFG guidance for conditional generation. During training, MODULI learns both a conditioned noise predictor $\epsilon_{\theta}(\bm{x}^{t},t,\bm{\omega},\bm{g})$ and an unconditioned one $\epsilon_{\theta}(\bm{x}^{t},t,\varnothing)$. We adopt a masking mechanism for training, with a probability of $10<p<$ 1 to zero out the condition of a trajectory, equivalent to an unconditional case. MODULI also employs a loss-weight trick, setting a higher weight for the next state $\bm{s}_{1}$ of the trajectory $\bm{x}^{0}(\tau)$ to encourage more focus on it, closely related to action execution.

### v2

@@ Abstract
Multi-objective Reinforcement Learning (MORL) seeks to develop policies that simultaneously optimize multiple conflicting objectives, but it requires extensive online interactions. Offline MORL provides a promising solution by training on pre-collected datasets to generalize to any preference upon deployment. However, real-world offline datasets are often conservatively and narrowly distributed, failing to comprehensively cover preferences, leading to the emergence of out-of-distribution (OOD) preference areas. Existing offline MORL algorithms exhibit poor generalization to OOD preferences, resulting in policies that do not align with preferences. Leveraging the excellent expressive and generalization capabilities of diffusion models, we propose MODULI (Multi Objective DiffUsion planner with sLIding guidance), which employs a preference-conditioned diffusion model as a planner to generate trajectories that align with various preferences and derive action for decision-making. To achieve accurate generation, MODULI introduces two return normalization methods under diverse preferences for refining guidance. To further enhance generalization to OOD preferences, MODULI proposes a novel sliding guidance mechanism, which involves training an additional slider adapter to capture the direction of preference changes. Incorporating the slider, it transitions from in-distribution (ID) preferences to generating OOD preferences, patching, and extending the incomplete Pareto front. Extensive experiments on the D4MORL benchmark demonstrate that our algorithm outperforms the state-of-the-art Offline MORL baselines, exhibiting excellent generalization to OOD preferences. Codes are available on this repository.
Generalization, Diffusion, Multi-Objective Reinforcement Learning, Slider Guidance, ICML
@@ 1 Introduction
Real-world decision-making tasks, such as robotic control (Liu et al., 2014; Cui et al., 2025; Yuan et al., 2025), autonomous driving (Li & Czarnecki, 2018), and industrial control (Bae et al., 2023), often require the optimization of multiple competing objectives simultaneously. It necessitates the trade-offs among different objectives to meet diverse preferences (Roijers et al., 2017; Yuan et al., 2024; Liu et al., 2024; Zhou et al., 2025). For instance, in robotic locomotion tasks, users typically focus on the robot’s movement speed and energy consumption (Xu et al., 2020). If the user has a high preference for speed, the agent will move quickly regardless of energy consumption; if the user aims to save energy, the agent should adjust to a lower speed. One effective approach is Multi-objective Reinforcement Learning (MORL), which enables agents to interact with vector-rewarded environments and learn policies that satisfy multiple preferences (Hayes et al., 2022; Roijers et al., 2013; Yang et al., 2019). These methods either construct a set of policies to approximate the Pareto front of optimal solutions (Alegre et al., 2023; Felten et al., 2024) or learn a preference-conditioned policy to adapt to any situation (Lu et al., 2023b; Hung et al., 2022). However, exploring and optimizing policies for each preference while alleviating conflicts among objectives, requires extensive online interactions (Alegre, 2023), thereby posing practical challenges due to high costs and potential safety risks.
@@
Offline MORL (Zhu et al., 2023a) proposes a promising solution that can learn preference-conditioned policies from pre-collected datasets with various preferences, improving data efficiency and minimizing interactions when deploying in high-risk environments. Most Offline MORL approaches focus on extending return-conditioned methods (Thomas et al., 2021; Zhu et al., 2023a) or encouraging consistent preferences through policy regularization (Lin et al., 2024). However, real-world offline datasets often come from users with different behavior policies, making it difficult to cover the full range of preferences and leading to missing preference regions, i.e., OOD preferences, resulting in a shattered or narrow data distribution. As shown in Figure 1, we visualize the trajectory returns for Hopper-Amateur dataset. The Complete version of the dataset can comprehensively cover a wide range of preferences. In many cases, suboptimal policies can lead to Shattered or Narrow datasets, resulting in OOD preference regions as marked by the red circles. We then visualize the approximate Pareto front for MORvS baseline (Zhu et al., 2023a) and our MODULI in such a shattered dataset. We find that the current Offline MORL baselines exhibit poor generalization performance when learning policies from incomplete datasets. They fail to adapt to OOD preference policies. The approximate Pareto front of MORvS shows significant gaps in the OOD preference areas, indicating that the derived policies and preferences are misaligned. Therefore, we aim to derive a general multi-objective policy that can model a wide range of preference conditions with sufficient expressiveness and strong generalization to OOD preferences. Motivated by the remarkable capabilities of diffusion models in expressiveness and generalization, we propose MODULI (Multi Objective DiffUsion planner with sLIding guidance), transforming the problem of Offline MORL into conditional generative planning. In Figure 1 (right), MODULI demonstrated a better approximation of the Pareto front, achieving returns beyond the dataset. MODULI also patched the missing preference regions and exhibited excellent generalization for the OOD preferences.
@@ 2 Related Work
Most MORL research primarily focuses on online settings, aiming to train a general preference-conditioned policy (Hung et al., 2022; Lu et al., 2023a) or a set of single-objective policies (Natarajan & Tadepalli, 2005; Xu et al., 2020) that can generalize to arbitrary preferences. This paper considers an offline MORL setting that learns policy from static datasets without online interactions. PEDI (Wu et al., 2021) combines the dual gradient ascent with pessimism to multi-objective tabular MDPs. MO-SPIBB (Thomas et al., 2021) achieves safe policy improvement within a predefined set of policies under constraints. These methods often require a priori knowledge about target preferences. PEDA (Zhu et al., 2023a) first develops D4MORL benchmarks for scalable Offline MORL with a general preference for high-dimensional MDPs with continuous states and actions. It also attempts to solve the Offline MORL via supervised learning, expanding return-conditioned sequential modeling to the MORL, such as MORvS (Emmons et al., 2021), and MODT (Chen et al., 2021). Lin et al. (2024) integrates policy-regularized into the MORL methods to alleviate the preference-inconsistent demonstration problem. This paper investigates the generalization problem of OOD preference, aiming to achieve better generalization through more expressive diffusion planners and the guidance of transition directions by a slider adapter.
@@
Diffusion models (DMs) (Ho et al., 2020; Song et al., 2021b) have become a mainstream class of generative models. Their remarkable capabilities in complex distribution modeling demonstrate outstanding performance across various domains (Kingma et al., 2021; Ruiz et al., 2023), inspiring research to apply DMs to decision-making tasks (Chi et al., 2023; Yang et al., 2024; Xian et al., 2023). Applying DMs for decision-making tasks can mainly be divided into three categories (Zhu et al., 2023b): generating long horizon trajectories like planner (Janner et al., 2022; Dong et al., 2024a; Ajay et al., 2023), serving as more expressive multimodal polices (Wang et al., 2023; Hansen-Estruch et al., 2023; Yuan et al., 2023) and performing as the synthesizers for data augmentation (Lu et al., 2024). Diffusion models exhibit excellent generalization and distribution matching abilities (Kadkhodaie et al., 2024; Li et al., 2024), and this property has also been applied to various types of decision-making tasks such as multi-task RL (He et al., 2024), meta RL (Ni et al., 2023) and aligning human feedback (Dong et al., 2023). In this paper, we further utilize sliding guidance to better stimulate the generalization capability of DMs, rather than merely fitting the distribution by data memorization (Gu et al., 2023).
@@ 3 Preliminaries
The MORL problem can be formulated as a multi-objective Markov decision process (MOMDP) (Chatterjee et al., 2006). A MOMDP with $n$ objectives can be represented by a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\Omega,\gamma\rangle$, where $\mathcal{S}$ and $\mathcal{A}$ demote the state and action spaces, $\mathcal{P}$ is the transition function, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{n}$ is the reward function that outputs $n$-dim vector rewards $\bm{r}=\mathcal{R}(\bm{s},\bm{a})$, $\Omega$ is the preference space containing vector preferences $\bm{\omega}$, and $\gamma$ is a discount factor. The policy $\pi$ is evaluated for $m$ distinct preferences $\{\bm{\omega}_{p}\}_{p=1}^{m}$, resulting policy set be represented as $\{\pi_{p}\}_{p=1}^{m}$, where $\pi_{p}=\pi(\bm{a}|\bm{s},\bm{\omega}_{p})$, and $\bm{G}^{\pi_{p}}=\mathbb{E}_{\bm{a}_{t}\sim\pi_{p}}\left[\sum_{t}\gamma^{t}\mathcal{R}(\bm{s}_{t},\bm{a}_{t})\right]$ is the corresponding unweighted expected return. We say the solution $\bm{G}^{\pi_{p}}$ is dominated by $\bm{G}^{\pi_{q}}$ when $G_{i}^{\pi_{p}}<G_{i}^{\pi_{q}}$ for $\forall i\in[1,2,\cdots,n]$. The Pareto front $P$ of the policy $\pi$ contains all its solutions that are not dominated. In MORL, we aim to define a policy such that its empirical Pareto front is a good approximation of the true one. While not knowing the true front, we can define a set of metrics for relative comparisons among algorithms, e.g., hypervolume, sparsity, and return deviation, to which we give a further explanation in Section 5.
@@
where $\bm{w}^{t}\in\mathbb{R}^{D}$ is the standard Wiener process, and $f(t)=\frac{{\rm d}\log\alpha_{t}}{{\rm d}t}$, $g^{2}(t)=\frac{{\rm d}\sigma^{2}_{t}}{{\rm d}t}-2\sigma^{2}_{t}\frac{{\rm d}\log\alpha_{t}}{{\rm d}t}$. The SDE forward process in Equation 1 has an probability flow ODE (PF-ODE) reverse process from time $T$ to 00 (Song et al., 2021b):
@@
Since $\bm{\epsilon}_{\theta}(\bm{x}^{t},t)$ can be considered as a predicted Gaussian noise added to $\bm{x}^{t}$, it is called the noise prediction model. With a well-train noise prediction model, DDIM (Song et al., 2021a) discretizes the PF-ODE to the first order for solving, i.e., for each sampling step:
@@
holds for any $0\leq t<s\leq T$. To enable the diffusion model to generate trajectories based on specified properties, Classifier Guidance (CG) (Dhariwal & Nichol, 2021) and Classifier-free Guidance (CFG) (Ho & Salimans, 2022) are the two main techniques. We use CFG due to its stable and easy-to-train properties. CFG directly uses a conditional noise predictor to guide the solver:
where $w$ is used to control the guidance strength.
Guided Sampling Methods
To enable the diffusion model to generate trajectories based on specified properties, Classifier Guidance (CG) (Dhariwal & Nichol, 2021) and Classifier-free Guidance (CFG) (Ho & Salimans, 2022) are the two main techniques. CG requires training an additional classifier $\log p_{\phi}(\bm{c}|\bm{x}^{t},t)$ to predict the log probability of exhibiting property $\bm{c}$. During inference, the gradient of the classifier is used to guide the solver:
$\bar{\bm{\epsilon}}_{\theta}(\bm{x}^{t},t,c)=\bm{\epsilon}_{\theta}(\bm{x}^{t},t)-w\cdot\sigma_{t}\nabla_{\bm{x}}\mathcal{C}_{\phi}(\bm{x}^{t},t,\bm{c}),$ (6)
on the contrary, CFG directly uses a conditional noise predictor to guide the solver:
@@ 4.1 Offline MORL as Conditional Generative Planning
We derive policies on the Offline MORL dataset $\mathcal{D}$ with trajectories $\tau=(\bm{\omega},\bm{s}_{0},\bm{a}_{0},\bm{r}_{0},\cdots,\bm{s}_{T-1},\bm{a}_{T-1},\bm{r}_{T-1})$ to respond to arbitrary target preferences $\bm{\omega}\in\Omega$ during the deployment phase. Previous Offline MORL (Zhu et al., 2023a) algorithms generally face poor expressiveness, unable to adapt to a wide range of preferences through a general preference-conditioned policy, resulting in sub-optimal approximate Pareto front. Inspired by the excellent expressiveness of diffusion models in decision making (Ajay et al., 2023; Dong et al., 2023), MODULI introduces diffusion models to model the policy. We define the sequential decision-making problem as a conditional generative objective $\max_{\theta}\mathbb{E}_{\tau\sim\mathcal{D}}\left[\log p_{\theta}\left(\boldsymbol{x}^{0}(\tau)\mid\boldsymbol{y}(\tau)\right)\right]$, where $\theta$ is the trainable parameter. The diffusion planning in MORL aims to generate sequential states $\bm{x}^{0}(\tau)=[\bm{s}_{0},\cdots,\bm{s}_{H-1}]$ with horizon $H$ to satisfy the denoising conditions $\bm{y}=[\bm{\omega},\bm{g}]$, where $\bm{\omega}$ is a specific preference. Similar to Zhu et al. (2023a), $\bm{g}_{t}$ is defined as the vector-valued Returns-to-Go (RTG) for a single step, i.e. $\bm{g}_{t}=\sum_{t}^{T}\mathcal{R}(\bm{s}_{t},\bm{a}_{t})$. The trajectory RTG $\bm{g}$ is the average of the RTGs at each step.
@@
Next, we can sample batches of $\bm{x}^{0}(\tau)$ and $\bm{y}$ from the dataset and update $p_{\theta}$ as the modified score matching objective in Equation 3:
$\mathcal{L}(\theta)=\mathbb{E}_{(\bm{x}^{0},\bm{\omega},\bm{g})\sim\mathcal{D},t\sim U(0,T),\epsilon\sim\mathcal{N}(\bm{0},\bm{I})}\left[\|\bm{\epsilon}_{\theta}(\bm{x}^{t},t,\bm{\omega},\bm{g})-\bm{\epsilon}\|^{2}_{2}\right].$
MODULI uses the CFG guidance for conditional generation. During training, MODULI learns both a conditioned noise predictor $\epsilon_{\theta}(\bm{x}^{t},t,\bm{\omega},\bm{g})$ and an unconditioned one $\epsilon_{\theta}(\bm{x}^{t},t,\varnothing)$. We adopt a masking mechanism for training, with a probability of $10<p<$ 1 to zero out the condition of a trajectory, equivalent to an unconditional case. We employ DiT1d (Dong et al., 2024b) network architecture instead of UNet (Ronneberger et al., 2015), which has been proven to have more precise generation capabilities. MODULI also employs a loss-weight trick, setting a higher weight for the next state $\bm{s}_{1}$ of the trajectory $\bm{x}^{0}(\tau)$ to encourage more focus on it, closely related to action execution. For details on the loss-weight trick, please refer to Appendix F.